OpenStudy (anonymous):

Can someone help me to check my answer of moment generating function of Y? it is in my last reply Let Y be a random variable with probability distribution. P(Y = y) = r*s^(y-1); for y = 1; 2; 3;...; and 0 < s < 1; Derive the value of r that makes this a valid probability distribution?And derive the moment generating function of Y?

OpenStudy (anonymous):

my answer is r = 1-s is this right?

OpenStudy (anonymous):

Hi @dumbcow, need your help again, you are my life saver!

OpenStudy (dumbcow):

hmm yes i believe you are right...that would make it a geometric distribution http://en.wikipedia.org/wiki/Geometric_distribution

OpenStudy (anonymous):

Thanks, the next step will be derive the moment generating function of Y. Hopefully I can do it by myself as well

OpenStudy (anonymous):

If it is a Prob Distr, it has to fulfill: \[\sum_{y=1}^{\infty}P(Y=y)=\sum_{y=1}^{\infty}r·s^{y-1}=1\rightarrow r=\frac{ 1 }{ \sum_{y=1}^{\infty}s^{y-1} } =\frac{ 1 }{ \frac{ 1 }{ 1-s}}=1-s\]

OpenStudy (anonymous):

Hi @dumbcow, Moment generating function of Y \[M(t)=E(e^{ty}) = \sum_{y=1}^{\infty} e^{ty} (1-s)s^{y-1} = (1-s) \sum_{y=1}^{\infty} e^{ty} s^{y-1} e^{t} e^{-t}\] \[= e^{t}(1-s)\sum_{y=1}^{\infty}e^{t(y-1)}s^{y-1} = e^{t}(1-s)\sum_{y=1}^{\infty}(e^{t}s)^{y-1} = \frac{ e^{t}(1-s) }{1-e^{t}s} \] where \[e^{t}s<1\rightarrow t<-\ln s\]